A car is stopped at a traffic light. It then travels along a straight road so that its distance from the light is given byπ‘₯(𝑑)=𝑏𝑑2βˆ’π‘π‘‘3,π‘€β„Žπ‘’π‘Ÿπ‘’π‘=2.4π‘šπ‘ βˆ’1π‘Žπ‘›π‘‘π‘=0.120π‘šπ‘ βˆ’3(a)Calculate the average velocity of the car for the time interval t = 0 to t = 10.0 s.(b)Calculate the instantaneous velocity of the car at t =0, t = 5.0 s and t = 10.0 s.


1
Expert's answer
2021-06-22T09:48:16-0400

"x (t) = bt^2-ct^3"

"a)"

"v_{avg}= \frac{x(t_2)-x(t_1)}{t_2-t_1};"

"x(0) = 0"

"x(10)= 2.4*10^2-0.12*10^3= 240-120 = 120"

"v_{avg}= \frac{x(10)-x(0)}{10- 0}= 12\frac{ m}{ s}"

"b)"

"v(t)= x'(t) = (bt^2-ct^3)'=2bt-3ct^2"

"v(0) =0"

"v(5)= 2*2.4*5-3*0.12*25=15\frac{ m}{ s}"

"v(10)=2*2.4*10-3*0.12*100=12\frac{ m}{ s}"

"\text{Answer :}a)v_{avg}= 12\frac{ m}{ s}"

"\text{ } b)v(0) =0;v(5)=15\frac{ m}{ s};v(10)=12\frac{ m}{ s}"